# Position paper: Do not explain (vision models) without context

### v2

@@ Abstract
Does the stethoscope in the picture make the adjacent person a doctor or a patient? This, of course, depends on the contextual relationship of the two objects. If it’s obvious, why don’t explanation methods for vision models use contextual information? In this paper, we (1) review the most popular methods of explaining computer vision models by pointing out that they do not take into account context information, (2) provide examples of real-world use cases where spatial context plays a significant role, (3) propose new research directions that may lead to better use of context information in explaining computer vision models, (4) argue that a change in approach to explanations is needed from where to how.
@@ 1 Introduction
The number of solutions incorporating Deep Learning (DL) models is rapidly increasing. Recently a significant change of the mindset in the community is observed. Many researchers agree that we should not only aim for the models with the highest performance but also the ones that behave in a responsible manner. This aspect is approached using eXplainable AI (XAI) techniques. In the work, we focus only on those for an investigation of Computer Vision models. The most attention is attributed to visual explanations. There are already many well-established post-hoc methods that generate colorful heatmaps highlighting the key regions in the input image for the model decision-making process. The major types of explanations are: permutation-based (i.e. LIME (Ribeiro et al., 2016)), gradient-based (i.e. Grad-CAM (Selvaraju et al., 2017), Integrated gradients (Sundararajan et al., 2017)) and propagation-based (i.e. Layer-wise Relevance Propagation (Binder et al., 2016), Concept Relevance Propagation (CRP) (Achtibat et al., 2022)). For a systematic survey on visual XAI methods, please consult the survey (Alicioglu & Sun, 2022; Holzinger et al., 2022).
The biggest attention of the XAI community is targeted towards the aforementioned visual explanations. In 2019, it was assessed that the number of papers on visual explanations was about 4 times bigger than the number of papers on textual and example-based explanations of models used in medical applications. The same trend was anticipated in 2020 (van der Velden et al., 2022). We suspect that this relationship has not changed much. The existing visual methods provide valuable insights, however, may not be enough. We identify a set of use cases when they would fail to explain accurately model’s predictions. These are the cases when the ground truth labels depend on the spatial relationships between objects constituting images. These may be observed in critical domains such as street surveillance systems for security, autonomous cars, and healthcare.
@@
The question is how to investigate whether the model takes into account spatial context. This area of model investigation seems largely unexplored. Recently, Achtibat et al. (2022) proposed to change the paradigm in XAI from where to what. It means that instead of simply highlighting the regions in the input images that are key for the model’s prediction, we should focus on extracting what semantic features within the images are important. We propose to shift the approach from where to how so that instead of only operating on the image pixel space where the pixels are highlighted, we should also analyse how the objects within images are oriented towards each other in the space (see Figure 1 for an illustration of the concept). The spatial relationships may be thought of as a kind of analogy to interaction terms that are investigated within models used for the analysis of tabular data.
As the field of XAI matures, we should focus on more complex concepts to capture the ambiguity of data and the intricate reasoning process of DL models. Note that the composition of the scene matters for humans when analysing images. This information can be captured using eye-trackers and the resulting saliency maps are widely studied and modelled using neural networks (Liu & Han, 2016). Therefore, investigating how may not only help in getting a deeper understanding of DL models, increasing safety by preventing inaccurate model predictions as we will have XAI tools to evaluate models before the launch but also in checking if there is an alignment between the way human and DL models reason about the scenes.
The structure of the work is as follows. First, we define the types of contextual information within images and provide illustrative examples (Section 2). Next, we elaborate on the domains when spatial context matters (Section 3). Then, we provide a summary of works on Deep Learning models where the spatial context plays a central part (Section 4). We contrast how much this field is developed with a survey on XAI methods that address the topic of spatial context (Section 6). Lastly, we outline the necessary research directions (Section 7) and draw conclusions (Section 8).
@@ 2 Contextual information within images
The key elements of scene compositions when analysing images using DL models were defined by Galleguillos & Belongie (2010): semantic context, spatial context and scale context. This nomenclature was built upon the taxonomy proposed in a psychological study (Biederman et al., 1982).
@@ 2.1 Semantic context
It refers to the situation when the co-presence of the objects in an image is helpful when performing a task i.e. classification.
@@ 2.1.2 Compound
We would like to add another category of semantic context that was not pointed out by Galleguillos & Belongie (2010): compound. There are objects i.e. rollerblades that are a compound of other self-existing objects (here, shoes and wheels). The information on whether there are wheels attached to shoes is crucial to classify the object correctly - shoes without wheels are not rollerblades anymore. Let us consider the possible outcome of the existing XAI methods when explaining the rollerblades prediction. When a model is only supposed to classify shoes vs rollerblades then it may be sufficient for the model to focus only on the detection of wheels and therefore this is the part that most probably will be highlighted within the explanation-heatmap. However, if there is a classification of rollerblades vs. cats then we would rather expect to see in the heatmap both the shoes and wheels highlighted. Therefore, we may face some relativity of explanations.
@@ 2.2.1 Spatial relationships
The fact that someone is standing within some distance from a shovel does not make this person a construction worker. The fact the man is holding a shovel increases the chances of him being a construction worker. Therefore, the distance is a key. It applies to different occupations and their attributes. Consider that someone is wearing a stethoscope (most probably a physician) and the other case when someone is touched only by a small part of the stethoscope during an examination (patient) (see Figure 6(d), 3(e)). Another example is already said rollerblades - not only there should be shoes and wheels in the picture but they should be properly attached. Similarly, an image can be classified as a bump if the cars are in contact otherwise if there is a significant distance between them, it is not a bump. This pairwise relationship can be extended to many objects i.e. when there is an image of some people spread far away from each other, we cannot classify the scene as a crowd unless people stand close to each other.
The distance can have also a more physics-informed aspect. Consider there is an image of a squared table where all the legs are attached to the same side of the table. Such a table would collapse even though it has the same elements as the typical table (table top and four legs). Note that even if the model will correctly distinguish that such a sample is out-of-distribution of the real-world tables, we would not be able to correctly explain the model’s decision using current XAI methods.
@@
Suppose there is a train near a tunnel. If the order of the objects is (1) railway engine, (2) carriage, (3) tunnel, it means that a train leaves the tunnel but if it is (1) carriage, (2) railway engine, (3) tunnel, then a train enters the tunnel (Figure 3(c), 3(g)). Also, when we think of a hamburger, we expect a piece of meat between two pieces of bread. If there are two pieces of bread and a piece of meat on top, this is not a proper hamburger.
@@
There are two cases regarding orientation: an angular position between objects (when we can assign some coordinate system to each object) and the situation when objects are heading or standing back towards each other. The example for the latter is a man giving roses to a woman (they face each other) vs. a man holding roses and standing back to a woman (Figure 3(d), 3(h)). It may seem a subtle case of order relationship because it can be translated to the question of whether the roses are between the two people or not, but we decide to analyse orientation as a separate subcategory. The other more severe case is whether a person with a knife/gun is heading towards the other person or is going in the other direction. In the first case, there is a high chance that the image depicts ‘violence’, whereas in the latter - it is more probable that there is ‘no violence’ in this particular frame.
@@ 2.2.2 Neighborhood
The spatial context can be also understood in a manner more similar to the way it is treated in NLP where it is analysed how many words surrounding the key part of the text we should input to the system. We would like to distinguish neighborhood as a separate subcategory of spatial context which can be thought of as co-presence but within a strictly constrained location (spatial aspect).
@@ 2.3 Scale context
It is difficult to assess the size of a given object if there is no reference object of a known size in the image. This may impact the correctness of a classification. Suppose a photo of a silver hoop - without contextual information, it can be either a ring or a bracelet.
@@ 3 Application domains
The videos from the streets and parks should be analysed in an automated way to detect danger and alarm the police. Automation is crucial as it would be difficult for people in municipal surveillance centres to analyse all of the recordings carefully in real time due to the abundance of data. The behavior of AI systems used for such an application should be trustworthy and safe (in compliance with regulations such as the EU AI Act and the Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence signed by President J. Biden). When analysing situations on streets, the case with a human holding a knife/gun in a relative position to another human is the scene that may occur. Therefore, the assessment of how severe the situations are, may depend on the orientation of the people towards each other.
@@
Safety is one of the biggest concerns in autonomous cars and so the same law regulations as in the case of surveillance systems apply. Suppose an autonomous car is approaching a crossing where a person is waiting to cross the street but there are no traffic signals. The vision system in an autonomous car identifies a human and a stick somewhere in the scene. Just based on this information, it is difficult to say whether the person is elderly and is holding a stick to facilitate walking or if it is elderly and a visible stick is just a part of a street sign or street lights. In the former, the car should predict that the needed time for the pedestrian to cross the street will be longer than in the latter and therefore, adjust its speed accordingly. Such adaptive cruise control can already benefit from the information about the surrounding cars to set the appropriate speed (Wei et al., 2020) and so should be done also in the case of approaching pedestrians.
@@
From the clinical point of view, it is important to check if lesions are clustered (primary) or widespread (with metastases) as it directly translates to the assessment of the stage of the disease and impacts the choice of treatment methods. For instance, in histopathology, the data is saved in the form of Whole Slide Images (WSIs) which are of huge resolution showing the whole cuff-out lesion at the level of individual cells. Therefore, the processing of them using Deep Learning can be troublesome. There are two possible approaches. First, analysis of only separate patches taken out of the whole WSI (if local labels are available), which leads to a loss of broader contextual information. Second, to analyse the whole lesion in one go to preserve spatial context (tissue-level perspective) (Lu et al., 2021; Shao et al., 2021; Chen et al., 2022; Rymarczyk et al., 2023; Lerousseau et al., 2021)
@@
Spatial context is important also in the case of synthetic data such as visual IQ tests in the form of Raven Progressive Matrices (RPM) (Raven & Court, 1998) where a test-taker is given a 3x3 grid of panels with one field missing (Figure 4(a) and the task is to choose the correct panel (matching the underlying relationship in the matrix) from the set of given panels (Figure 4(b)). The relationships within the images in the panels can occur within rows, columns and/or diagonals of the matrix. An example of the underlying rule within the RPM can be that there is a diamond shape in each panel that is rotated by the same angle from one panel to another one in each row but the starting orientations of the diamonds in the first column are different. Therefore, the RPM has to be analysed as a whole, paying attention to spatial relationships between the objects in panels. For throughout survey on the application of Deep learning to abstract reasoning tasks (including visual IQ tests) consult a work of Małkiński & Mańdziuk (2023).
@@ 4 Spatial context in DL models
Spatial context can be valuable when performing Deep Learning tasks. We analyse whether the information on spatial aspects is concerned within the design (architectures) and training schemes of DL models.
@@
In the early days of neural networks, the images were flattened before being passed to multilayer perceptrons, which led to an irrevocable loss of spatial aspect within the data. In order to overcome this downside, convolutional neural networks were proposed that analyse the images as matrices. Such a design of convolutional layers was supposed to capture local patterns. Later, a new architecture (Vision Transformer) was introduced with a self-attention mechanism as a main part. The self-attention is supposed to retrieve (even the long-range) relationships between the input patches. Additionally, the model is fed with patch positional encodings.
Within this timeline of consistent progress in the field, two solutions outside of the main track deserve attention: CapsuleNet (Sabour et al., 2017), MLP-Mixer (Tolstikhin et al., 2021). The motivation behind the idea of CapsuleNets was to propose an architecture that will take into account the spatial relationships between objects in the image as opposed to convolutional neural networks (CNNs) that were claimed to pay attention only to different elements without spatial context due to continuous downsampling operations (Pawan & Rajan, 2022). Such a by-design behaviour of CNNs was perceived as a drawback. The illustrative example is provided in Figure 5.
Another worth-mentioning solution that went off the beaten track was MLP-Mixer. It was introduced before the Vision Transformer. The authors regardless of the trend to benefit from contextual information, showed in one of the experiments, that MLP-Mixer is invariant to the order of input patches and pixels within the patches treating it as an asset. The performance of the MLP-Mixer was the same on the said modified images as on the original ones. In the case of ResNet, the performance on the images where the spatial relationships were corrupted in the said manner, dropped significantly. It is claimed that it is due to CNNs’ strong inductive bias.
@@
The topic was analysed for natural scene images (mostly in a detection task). The early approaches are summarized in a survey (Galleguillos & Belongie, 2010). The vast number of works incorporated the concept of neighborhood into the pipeline. One of those is spatial-context-aware deep neural network (Zhang et al., 2022) used for multi-label classification when bounding boxes are provided. In the solution, there are two bounding box generation branches - the object and context ones. The first one localizes precisely the objects (tight bounding boxes) whereas the second expands the bounding boxes by incorporating the neighborhood around the object. Later the features are extracted from the two branches and combined to perform a final classification benefitting from contextual information.
@@
Histopathological data: The concept of incorporating information from various resolutions was a central part of several works (Chen et al., 2022; Zormpas-Petridis et al., 2019; Marini et al., 2021). Another concept was to use information from neighboring patches of the one that was under analysis at a particular moment. Zhang et al. (2020) propose two 2D-LSTM spatial context branches on top of a CNN feature extractor to obtain continuous and discrete spatial context-dependent information.
@@
The spatial aspect was also incorporated into model pretraining schemes. One class of such approaches are the ones dedicated to inpainting where the model is asked to fill up the missing parts of the input image. In such cases, the generation of missing image parts is conditioned on its surroundings (i.e. Pathak et al., 2016; He et al., 2022). The resulting context encoder can be later used for various downstream tasks as they capture the semantics of visual structures.
The pretraining scheme where the idea of spatial context was incorporated in a more explicit form is described by Doersch et al. (2015). The model is given two neighboring patches of the image and is asked to output the spatial configuration of them out of 8 possibilities. There is a grid 3x3 where the first image is set and the position of the second within the grid is the target.
@@
In all of the aforementioned works except the one by Doersch et al. (2015), the understanding of spatial context by DL models is not the goal itself, it is a means towards better performance in the main task. In the work of Swingler & Bath (2020), the prediction of spatial context is the centre of the experiments. In the first experiment, the model is given an image of two cloth pieces and it is supposed to classify each object and classify what type of spatial relationship is present (left, above, below left, and below right). In the second experiment, one sample input is an image containing three objects and one-hot embeddings of two objects. The task is to output the spatial relation between the two objects mentioned within the input. The focus on spatial relationships within the output is also in generative models when relative positions of the objects are specified in the prompts (Wang et al., 2024).

### v3

# Position: Do Not Explain Vision Models Without Context

@@ Abstract
Does the stethoscope in the picture make the adjacent person a doctor or a patient? This, of course, depends on the contextual relationship of the two objects. If it’s obvious, why don’t explanation methods for vision models use contextual information? In this paper, we (1) review the most popular methods of explaining computer vision models by pointing out that they do not take into account context information, (2) show examples of failures of popular XAI methods, (3) provide examples of real-world use cases where spatial context plays a significant role, (4) propose new research directions that may lead to better use of context information in explaining computer vision models, (5) argue that a change in approach to explanations is needed from where to how.
@@ 1 Introduction
The number of solutions that incorporate Deep Learning (DL) models is rapidly increasing. Recently, a significant change of the mindset in the community is observed. Many researchers agree that we should not only aim for the models with the highest performance but also the ones that behave in a responsible manner. This aspect is approached using eXplainable AI (XAI) techniques. In the work, we focus only on those for an investigation of Computer Vision models. The most attention is attributed to visual explanations. There are already many well-established post-hoc methods that generate colorful heatmaps highlighting the key regions in the input image for the model decision-making process. The major types of explanations are: permutation-based (i.e. LIME (Ribeiro et al., 2016)), gradient-based (i.e. Grad-CAM (Selvaraju et al., 2017), Integrated gradients (Sundararajan et al., 2017)) and propagation-based (i.e. Layer-wise Relevance Propagation (Binder et al., 2016), Concept Relevance Propagation (Achtibat et al., 2022)). For a systematic survey on visual XAI methods, please consult a survey (Alicioglu & Sun, 2022; Holzinger et al., 2022).
The biggest attention of the XAI community is focused on the aforementioned visual explanations. In 2019, it was assessed that the number of papers on visual explanations was about 4 times bigger than the number of papers on textual and example-based explanations of models used in medical applications. The same trend was anticipated in 2020 (van der Velden et al., 2022). We suspect that this has not changed much. The existing visual methods provide valuable insights, however, may not be enough. We identify a set of use cases in which they would fail to accurately explain the model’s predictions. These are the cases where the ground truth labels depend on the spatial relationships between objects constituting images. These may be observed in critical domains, such as street surveillance systems, autonomous cars, and healthcare. This position paper argues that we should take into account contextual information when explaining vision models – we need spatial XAI.
@@
The question is how to investigate whether the model takes into account spatial context. This area of model investigation seems largely unexplored. Recently, Achtibat et al. (2022) proposed to change the paradigm in XAI from where to what. It means that instead of simply highlighting the regions in the input images that are key for the model’s prediction, we should focus on extracting what semantic features within the images are important. We propose to shift the approach from where to how so that instead of only operating on the image pixel space where the pixels are highlighted, we should also analyze how the objects within images are oriented towards each other in the space (Figure 1). The spatial relationships may be thought of as a kind of analogy to interaction terms that are investigated in models for tabular data analysis.
As the field of XAI matures, we should focus on more complex concepts to capture the ambiguity of data and the intricate reasoning of DL models. Humans when analyzing images often take into account the composition of the scene. This information can be captured using eye trackers. The resulting saliency maps are widely studied and modeled using neural networks (Liu & Han, 2016). Therefore, investigating how may not only help in getting a deeper understanding of DL models, increasing safety by preventing inaccurate model predictions as we will have XAI tools to evaluate models before the launch but also in checking if there is an alignment between the way human and DL models reason about the scenes.
The structure of the work is as follows. First, we define the types of contextual information within images and provide illustrative examples (Section 2). Next, we elaborate on the domains when spatial context matters (Section 3). Then, we provide a summary of works on DL models where the spatial context plays a central role (Section 4). We provide examples of failures of popular XAI techniques in Section 5. Later, we contrast the number solutions addressing the topic of spatial context developed in the modelling community with the number of papers in the topic of spatial XAI (Section 6). Lastly, we outline the necessary research directions (Section 7) and draw conclusions (Section 8).
@@ 2 Contextual Information within Images
The key elements of scene compositions when analyzing images using DL models were defined by Galleguillos & Belongie (2010): semantic context, spatial context and scale context. This nomenclature was built upon the taxonomy proposed in a psychological study (Biederman et al., 1982).
@@ 2.1 Semantic Context
It refers to the situation when the co-presence of the objects in an image is helpful when performing a task.
@@ 2.1.2 Compound
We would like to add another category of semantic context that was not pointed out by Galleguillos & Belongie (2010): compound. There are objects i.e. rollerblades that are a compound of other self-existing objects (here, shoes and wheels). The information on whether there are wheels attached to shoes is crucial to classify the object correctly – shoes without wheels are not rollerblades anymore. Let us consider the possible outcome of the existing XAI methods when explaining the rollerblades prediction. When a model is only supposed to classify shoes vs rollerblades then it may be sufficient for the model to focus only on the detection of wheels and therefore this is the part that most probably will be highlighted within the explanation heatmap. However, if there is a classification of rollerblades vs. cats then we would rather expect to see in the heatmap both the shoes and wheels highlighted. Therefore, we may face some relativity of explanations.
@@ 2.2.1 Spatial relationships
The fact that someone is standing within some distance from a shovel does not make this person a construction worker. The fact the man is holding a shovel increases the chances of him being a construction worker. Therefore, the distance is a key. It applies to different occupations and their attributes. Consider that someone is wearing a stethoscope (most probably a physician) and the other case when someone is touched only by a small part of the stethoscope during an examination (patient) (see Figure 3(a), 3(e)). Another example is already said rollerblades – not only there should be shoes and wheels in the picture but they should be properly attached. Similarly, an image can be classified as a bump if the cars are in contact otherwise if there is a significant distance between them, it is not a bump. This pairwise relationship can be extended to many objects i.e. when there is an image of some people spread far away from each other, we cannot classify the scene as a crowd unless people are standing close to each other.
The distance can have also a more physics-informed aspect. Consider there is an image of a squared table where all the legs are attached to the same side of a table. Such a table would collapse even though it has the same elements as the typical table (table top and four legs). Note that even if the model will correctly distinguish that such a sample is outside of a distribution of the real-world tables, we would not be able to correctly explain the model’s decision using current XAI methods.
@@
Suppose there is a train near a tunnel. If the order of the objects is (1) railway engine, (2) carriage, (3) tunnel, it means that a train is leaving the tunnel but if it is (1) carriage, (2) railway engine, (3) tunnel, then a train is entering the tunnel (Figure 3(c), 3(g)). Also, when we think of a hamburger, we expect a piece of meat between two pieces of bread. If there are two pieces of bread and a piece of meat on top, this is not a proper hamburger.
@@
There are two cases regarding orientation: an angular position between objects (when we assign some coordinate system to each object) and the situation when objects are heading or standing back towards each other. The example for the latter is a man giving roses to a woman (they face each other) vs. a man holding roses but standing back to a woman (Figure 3(d), 3(h)). It may seem a subtle case of order relationship because it can be translated to the question of whether the roses are between the two people or not, but we decide to treat it as a sub-case of orientation. The other more severe case is whether a person with a knife/gun is heading towards the other person or is going in the other direction. In the first case, there is a high chance that the image depicts ‘violence’, whereas in the latter – it is more probable that there is ‘no violence’ in this particular frame.
@@ 2.2.2 Neighborhood
The spatial context can be also understood in a manner more similar to the way it is treated in NLP where it is analyzed how many words surrounding the key part of the text we should input to the system. We would like to distinguish neighborhood as a separate subcategory of spatial context which can be thought of as co-presence but within a strictly constrained location (spatial aspect).
@@ 2.3 Scale Context
It is difficult to assess the size of a given object if there is no reference object of a known size in the image. This may impact the correctness of a classification. Suppose a photo of a silver hoop – without contextual information, it can be either a ring or a bracelet.
@@ 3 Application Domains
The videos from the streets and parks should be analyzed in an automated way to detect danger and alarm the police. The behavior of such AI systems should be trustworthy and safe (in compliance with regulations such as the EU AI Act and the Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence signed by President J. Biden). When analyzing situations on streets, the case with a human holding a knife/gun in a relative position to another human may occur. Therefore, the assessment of how severe the situations are, may depend on the orientation of the people towards each other.
@@
Safety is one of the biggest concerns in autonomous cars and so the same law regulations as in the case of surveillance systems apply. Suppose an autonomous car is approaching a crossing where a person is waiting to cross the street but there are no traffic signals. The vision system in an autonomous car identifies a human and a stick somewhere in the scene. Just based on this information, it is difficult to say whether the person is elderly and is holding a stick to facilitate walking or if it is elderly and a visible stick is just a part of a street sign or street lights. In the former, the car should predict that the needed time for the pedestrian to cross the street will be longer than in the latter and therefore, adjust its speed accordingly. Such adaptive cruise control can already benefit from the information about the surrounding cars to set the appropriate speed (Wei et al., 2020) and so should be done also in the case of pedestrians.
@@
From the clinical point of view, it is important to check if lesions are clustered (primary) or widespread (with metastases) as it directly translates to the assessment of the stage of the disease and impacts the choice of treatment methods. For instance, in histopathology, the data is saved in the form of Whole Slide Images (WSIs) which are of huge resolution showing the whole cuff-out lesion at the level of individual cells. Therefore, the processing of them using Deep Learning can be troublesome. There are two possible approaches: (1) analyze only separate patches taken out of the whole WSI (if local labels are available), which leads to a loss of broader contextual information, (2) analyze the whole lesion in one go to preserve spatial context (tissue-level perspective) (Lu et al., 2021; Shao et al., 2021; Chen et al., 2022; Rymarczyk et al., 2023).
@@
Spatial context is important also in the case of synthetic data such as visual IQ tests in the form of Raven Progressive Matrices (RPM) (Raven & Court, 1998) where a test-taker is given a grid of panels with one field missing (Figure 4(a)) and the task is to choose the correct panel (matching the underlying relationship in the matrix) from the set of given panels (Figure 4(b)). The relationships within the images in the panels can occur within rows, columns and/or diagonals of the matrix. An example of the underlying rule within the RPM can be that there is a diamond shape in each panel that is rotated by the same angle from one panel to another one in each row but the starting orientations of the diamonds in the first column are different. Therefore, the RPM has to be analyzed as a whole, paying attention to spatial relationships between the objects in panels. For a throughout survey on the application of Deep Learning to abstract reasoning tasks (including visual IQ tests) consult a work of Małkiński & Mańdziuk (2023).
@@ 4 Spatial Context in DL Models
Spatial context can be valuable when performing Deep Learning tasks. We analyze whether the information on spatial aspects is concerned within the design (architectures) and training schemes of DL models.
@@
In the early days of neural networks, the images were flattened before being passed to multilayer perceptrons, which led to an irrevocable loss of spatial aspect within the data. In order to overcome this downside, convolutional neural networks (CNNs) were proposed that analyze the images as matrices. Such a design of convolutional layers was supposed to capture local patterns. Later, a new architecture, Vision Transformer, was introduced with a self-attention mechanism as a main part. The self-attention is supposed to retrieve even the long-range relationships between the input patches. Additionally, the model is fed with patch positional encodings.
Within this timeline of consistent progress in the field, two solutions outside of the main track deserve attention: CapsuleNet (Sabour et al., 2017), MLP-Mixer (Tolstikhin et al., 2021). The motivation behind the idea of CapsuleNets was to propose an architecture that will take into account the spatial relationships between objects in the image as opposed to CNNs that were claimed to pay attention only to different elements of the image without considering spatial context due to continuous downsampling operations (Pawan & Rajan, 2022). Such a by-design behavior of CNNs was perceived as a drawback. The illustrative example is provided in Figure 5.
Another worth-mentioning solution that went off the beaten track was MLP-Mixer. The authors regardless of the trend to benefit from contextual information, showed in one of the experiments, that MLP-Mixer is invariant to the order of input patches and pixels within the patches treating it as an asset. The performance of the MLP-Mixer was the same on the said modified images as on the original ones. In the case of ResNet, the performance on the images where the spatial relationships were corrupted in the said manner, dropped significantly. It is claimed that it is due to CNNs’ strong inductive bias.
@@
The topic was analyzed for natural scene images (mostly in a detection task). The early approaches are summarized in a survey (Galleguillos & Belongie, 2010). The vast number of works incorporated the concept of neighborhood into the pipeline. One of those is spatial-context-aware deep neural network (Zhang et al., 2022) used for multi-label classification when bounding boxes are provided. In the solution, there are two bounding box generation branches – the object and context ones. The first one localizes precisely the objects (tight bounding boxes) whereas the second expands the bounding boxes by incorporating the neighborhood around the object. Later the features are extracted from the two branches and combined to perform a final classification benefitting from contextual information.
@@
Histopathological data: The concept of incorporating information from various resolutions was a central part of several works (Chen et al., 2022; Zormpas-Petridis et al., 2019; Marini et al., 2021). Another approach was to use information from neighboring patches of the one that was under analysis at a particular moment. Zhang et al. (2020) propose two 2D-LSTM spatial context branches on top of a CNN feature extractor to obtain continuous and discrete spatial context-dependent information.
@@
The spatial aspect was also incorporated into model pretraining schemes. One class of such approaches is inpainting where the model is asked to fill up the missing parts of the input image. In such cases, the generation of missing image parts is conditioned on its surroundings (i.e. Pathak et al., 2016; He et al., 2022). The resulting context encoder can be later used for various downstream tasks as they capture the semantics of visual structures.
The pretraining scheme where the idea of spatial context was incorporated in a more explicit form is described by Doersch et al. (2015). The model is given two neighboring patches from an image and is asked to output their spatial configuration out of 8 possibilities. There is a grid where the first patch is set and the position of the second within the grid is a target.
@@
In all of the aforementioned works except the one by Doersch et al. (2015), the understanding of spatial context by DL models is not the goal itself but a means towards better performance in the main task. In the work of Swingler & Bath (2020), the prediction of spatial context is the center of the experiments. In the first experiment, the model is given an image of two cloth pieces and is supposed to classify each object and output the type of spatial relationship that is present (left, above, below left, and below right). In the second experiment, one sample input is an image containing three objects and one-hot embeddings of two objects. The task is to output the spatial relation between the two objects mentioned within the input embeddings.
The spatial relationships within the output are also investigated in generative models when relative positions of the objects are specified in the prompts (Wang et al., 2024).
@@
We acknowledge that there are some recent solutions dedicated to scene understanding i.e. VisProg (Gupta & Kembhavi, 2023), ViperGPT (Surís et al., 2023) and physics-informed NNs (Raissi et al., 2019) that potentially could understand that the right position of the legs in the table prevents collapse. This shows that the topic of spatial context is recognized as important in the community working on DL models. However, we show that it is rather overlooked in the XAI community. These recent DL solutions are not strictly vision models, which are of the biggest interest in the XAI community where the focus is on investigation of the mature solutions like CNNs and ViTs that already find real-life applications.